Determine the location and values of the absolute maximum and absoluteย 

minimum for the given function:ย 

F(x) = (โˆ’x + 2)

Power 4

, ๐‘คโ„Ž๐‘’๐‘Ÿ๐‘’ 0 โ‰ค x โ‰ค 3


1
Expert's answer
2021-06-14T13:57:39-0400
"F(x)=(-x+2)^4"

"0\leq x\leq 3"

Find the first derivative


"F'(x)=((-x+2)^4)'=-4(-x+2)^3"

Find the critical number(s)


"F'(x)=0=>-4(-x+2)^3=0"

"x=2"

Critical number: "2"


"F(0)=(-2+0)^4=16"


"F(3)=(-2+3)^4=1"

"F(2)=(-2+2)^4=0"

The function "F(x)" has the absolute maximum with value of "16" on "[0, 3]" at "x=0: Point(0,16)."


The function "F(x)" has the absolute minimum with value of on "[0, 3]" at "x=2: Point(2,0)."